OpenStudy (anonymous):

Describe a real world scenario that is an arithmetic series. Show your work as you find the 7th term of your scenario. HELP!!!

OpenStudy (campbell_st):

a stack of cut logs the bottom row has 23 logs and each row above has one less log  how many logs in the 7th row a = 23 d = 1 and n = 7 use the formula \[T _{n} = a + (n -1) d\]

OpenStudy (anonymous):

@campbell_st can you solve it for me please

OpenStudy (campbell_st):

oops in the information above the common difference is d = -1 \[T _{7} = 23 + (7 -1)\times (-1)\] \[T _{7} = 23 - 6\] I'll let you finish it

OpenStudy (anonymous):

@campbell_st so, its 17???

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

@campbell_st thanks I was having trouble thinjking of a real world example the rest is fairly easy. Thanks A LOT

OpenStudy (campbell_st):

hope it helps... things like seats in a theatre is another example...